The difference of the two numbers is 11. The second is 8 less than 2 times the first. What are the two numbers?
Answer by stanbon(75887)

The difference of the two numbers is 11. The second is 8 less than 2 times the first. What are the two numbers?
--------
Let the numbers be "x" and "y".
EQUATIONS:
1st: x-y=11 (this means the x is larger than the y)
2nd: y=2x-8
-----
Substitute 2nd into 1st to solve for "x":
x-(2x-8) = 11
x-2x+8 = 11
-x=3
x=-3
-------
Substitute that into y=2x-8 to solve for "y":
y = (2*-3) - 8
y = -14
-------------
Cheers,
Stan H.